# Predicting wide binaries and deviations from standard gravity using machine learning algorithms

### v2

@@ V Conclusions
We foresee several interesting directions that can be taken up as future problems, such as predicting deviations from the raw GAIA dataset and not the El-Badry catalogue, integrating and merging the repository on WBS predictions and anomaly detection on the raw GAIA data, expanding the models to predict more exotic and general gravitational phenomena or building an ML-based stellar object identifier for the GAIA data.
@@ A.2 Detailed plots and tables
Appendix B Predicting anomalies using ML
B.1 Data characteristics
The class distribution in the filtered dataset was:
fdt[’target’].value_counts()
    0    80350
    1      738
train[’target’].value_counts()
    0    64280
    1      590
test[’target’].value_counts()
    0    16070
    1      148
B.2 Detailed plots and tables
Table 10: Performance Comparison of the most accurate models (trained on correlation cutoff dataset)
Algorithms Precision Recall F1 score Accuracy
RFC 0.838710 0.527027 0.647303 0.994759
LR 0.000000 0.000000 0.000000 0.990874
SVM (RBF) 0.000000 0.000000 0.000000 0.990874
DTC 0.496933 0.547297 0.520900 0.990813
AdaBoost 0.752475 0.513514 0.610442 0.994019
KNN 0.000000 0.000000 0.000000 0.990874
NB 0.000000 0.000000 0.000000 0.990874
Bagging 0.824176 0.506757 0.627615 0.994512
Table 11: Classification Metrics of the most accurate algorithms trained on the correlation cutoff dataset
RFC 78 52.702703 85 57.432432
LR 0 0.000000 148 100.000000
SVM (RBF) 0 0.000000 148 100.000000
DTC 81 54.729730 149 100.675676
AdaBoost 76 51.351351 97 65.540541
KNN 0 0.000000 148 100.000000
NB 0 0.000000 148 100.000000
Bagging 75 50.675676 89 60.135135

### v3

@@ 4 Conclusions
As part of an ongoing work, we plan to extend the scope of application of machine learning in the context of wide binaries to identify potential anomalous wide binaries by casting it as a supervised anomaly detection problem. This task would be a step towards using ML models to learn the characteristics of systematic deviations from newtonian gravity in a population sample of wide binaries as reported in recent literature [Hernandez et al.(2021)Hernandez, Cookson, and Cortés, Hernandez et al.(2024)Hernandez, Chae, and Aguayo-Ortiz, Hernandez et al.(2012)Hernandez, Jiménez, and Allen, Hernandez(2023), Hernandez et al.(2023)Hernandez, Verteletskyi, Nasser, and Aguayo-Ortiz, Chae(2023a), Chae(2024), Chae(2022), Chae(2023b)]. We foresee several interesting directions that can be taken up as future problems, such as integrating and merging the repository on WBS predictions and anomaly detection on the raw Gaia data, expanding the models to predict more exotic and general gravitational phenomena or building an ML-based stellar object identifier for the Gaia data.